OpenStudy (anonymous):

One solution of the equation 4x2 - 9x + 5 = 0 is five over four . What is the other solution? Answer

hero (hero):

4x^2 - 4x - 5x + 5 = 0 4x(x-1) -5(x-1) = 0 (x-1)(4x-5) = 0 The other solution is x = 1

OpenStudy (anonymous):

that means that one bracket is ( 4x - 5) so the other must b (x - a) - can u see what a must be?

OpenStudy (anonymous):

Trick: Sum of the coefficients is 0, hence 1 is a solution ;-)

OpenStudy (anonymous):

good trick